### v1

# Talking to DINO: Bridging Self-Supervised Vision Backbones with Language for Open-Vocabulary Segmentation

Abstract
Open-Vocabulary Segmentation (OVS) aims at segmenting images from free-form textual concepts without predefined training classes. While existing vision-language models such as CLIP can generate segmentation masks by leveraging coarse spatial information from Vision Transformers, they face challenges in spatial localization due to their global alignment of image and text features. Conversely, self-supervised visual models like DINO excel in fine-grained visual encoding but lack integration with language. To bridge this gap, we present Talk2DINO, a novel hybrid approach that combines the spatial accuracy of DINOv2 with the language understanding of CLIP. Our approach aligns the textual embeddings of CLIP to the patch-level features of DINOv2 through a learned mapping function without the need to fine-tune the underlying backbones. At training time, we exploit the attention maps of DINOv2 to selectively align local visual patches with textual embeddings. We show that the powerful semantic and localization abilities of Talk2DINO can enhance the segmentation process, resulting in more natural and less noisy segmentations, and that our approach can also effectively distinguish foreground objects from the background. Experimental results demonstrate that Talk2DINO achieves state-of-the-art performance across several unsupervised OVS benchmarks. Source code and models are publicly available at: https://lorebianchi98.github.io/Talk2DINO/. Equal contribution.
1 Introduction
Figure 1: Overview. Our approach aligns the embedding spaces of CLIP and DINOv2 through a learned mapping function. This results in fine-grained visual encodings, which can be mapped to language to provide natural and less noisy semantic segmentation.
Open-Vocabulary Segmentation (OVS) [54] is a fundamental task in Computer Vision that aims to partition an input image into a set of coherent regions based on concepts provided at inference time [18, 24, 6]. The set of concepts used to partition the image is usually provided in freeform natural language, which effectively unchains the methods from working with a specific fixed set of classes provided at training time. Solving this task requires a fine-grained understanding of the semantic interconnections between image pixels and the meaning conveyed by natural language.
While previous works employed pixel-level annotations as a source of supervision [48, 53, 60], a recent trend in the field is to tackle this problem in an unsupervised manner [49, 9, 3] leveraging the correlations learned by state-of-the-art backbones. Contrastive embedding spaces like CLIP [34], in particular, have demonstrated good performance on tasks that demand a holistic understanding of vision and language modalities [55, 25, 29, 59], and have therefore been employed for unsupervised OVS [58, 42, 23]. Although CLIP-based backbones exhibit strong cross-modal capabilities, they are primarily trained to predict a global similarity score between text and images, which limits their spatial understanding and consequently affects tasks based on dense predictions. Recent efforts have tackled this limitation by introducing architectural modifications [58, 42, 19]. However, the spatial understanding constraints imposed by the training modality hinder the effectiveness of such backbones in OVS and highlight the potential benefits of exploring alternative models with enhanced perceptual capabilities.
Self-supervised vision-only backbones like DINO and DINOv2 [8, 33, 12] have instead shown remarkable abilities in capturing fine-grained and localized spatial features without the reliance on annotated data. Specifically, the self-attention mechanism in such backbones generates attention maps that consistently pinpoint relevant regions within the image and has been widely leveraged for foreground object segmentation [38, 39, 45, 44, 43]. While this property makes them a powerful tool for tasks requiring fine-grained spatial understanding, the embedding space derived from visual self-supervised networks is not inherently aligned with textual concepts, making it incompatible with the OVS task.
To close the gap between vision-language and self-supervised embedding spaces, we propose Talk2DINO, a method that combines the spatial sensitivity of DINOv2 with the text-image alignment capabilities of CLIP, enabling a highly localized, multimodal image understanding. Our approach, depicted in Fig. 1, learns a mapping function that translates the text embeddings of CLIP to interact with DINOv2’s patch-level embeddings without fine-tuning the underlying backbones. In particular, our alignment mechanism enhances text-image correspondence by exploiting self-attention heads of DINOv2 to highlight diverse semantic regions of the image. During training, we weight visual patch features using DINOv2’s attention maps, dynamically selecting the head that best aligns with the provided caption. This visual feature is then used to maximize similarity with the caption through contrastive learning. At inference, we calculate the similarity of visual patches to each textual label, including a novel background cleaning procedure that re-weights class scores based on the self-attention maps.
Our approach demonstrates state-of-the-art performance on several unsupervised OVS datasets by only learning a minimal set of parameters. Our results highlight that a self-supervised vision-only encoder can generate embeddings with semantic properties akin to textual representations, opening up new pathways for addressing spatial understanding limitations in CLIP-like models [40].
To sum up, our main contributions are as follows:
  • We propose Talk2DINO, the first model that directly aligns DINOv2 and CLIP feature spaces for OVS. By mapping CLIP’s textual embeddings into the DINOv2 space using a non-linear warping function, Talk2DINO effectively provides language properties to DINOv2.
  • Our proposed model employs a novel training schema that selects the most relevant visual self-attention head and does not need fine-tuning on the backbones.
  • We showcase the capabilities of Talk2DINO on unsupervised OVS by devising a computationally efficient inference pipeline that also employs a novel approach based on DINOv2 self-attention to clean background objects.
  • Experimentally, we show that Talk2DINO achieves state-of-the-art results in standard OVS benchmarks, demonstrating the effectiveness of the proposed approach.
2 Related Work
Vision-Language Pre-Training. In the last years, vision-language pre-training has gained increasing interest by learning multimodal representations that can be easily transferred to downstream tasks [16, 34, 21]. The popular CLIP model [34] is trained on large-scale web-scraped data through a contrastive objective, which maximizes the similarity of the representations of the corresponding image-text pair while minimizing the similarity of the other pairs within a batch. This approach has demonstrated remarkable zero-shot classification and retrieval performance. However, learning a multimodal representation by matching global images and texts poses challenges in localizing regions with their corresponding text, showing limited performance in tasks based on dense predictions [31, 35, 55, 4].
Open-Vocabulary Segmentation. In zero-shot segmentation, a segmentation model is trained on a set of seen classes and must generalize to unseen classes. The first attempts of OVS inherit this paradigm by training the model on a closed set of classes for which segmentation data is available and exploiting vision-language pre-training to extend their capabilities on an open set of classes through text [53, 10, 48, 20, 60]. The two-step approach represents the most popular framework in this research direction, which proposes class-agnostic regions, trained on segmentation masks, and provides them to CLIP to be aligned with textual classes [13, 18, 26, 52, 51]. However, this approach is affected by performance gaps between seen and unseen classes and presents a significant computational overhead.
On the contrary, another research direction investigates how to force the segmentation capabilities to emerge without relying on direct supervision from segmentation data. Cha et al. [9] identify this setting as unsupervised OVS and propose a unified evaluation protocol. Approaches in this line can be categorized into two main groups: (i) training-free methods that propose architectural adaptations to enable pre-trained models to produce localized multimodal features [58, 23, 46, 5, 41, 42, 19], and (ii) weakly-supervised methods that leverage large sets of image-text pairs with dedicated learning strategies that aim to improve the correspondence between regions and texts [49, 9, 35, 50, 28, 36]. Our model lies in the former category since we exploit a weak supervision to learn how to bridge the CLIP and DINOv2 feature spaces.
Self-Supervised Backbones. Recent advances in self-supervised learning have led to models that showcase impressive matching and localization capabilities. In particular, the DINO family of models [8, 33, 12] employs Vision Transformers [14] trained with a self-distillation loss function, and has shown that patch-level features learned through self-supervision can yield semantic information. Moreover, a strong relationship has been observed between self-attention activations and foreground regions of the image. Hence, researchers have recently focused on the usage of DINO for the unsupervised object segmentation task [38, 39, 45, 44, 43].
While the aforementioned models are purely visual, recent works have also focused on connecting self-supervised feature spaces with textual representations for OVS. ReCo [37], OVDiff [22], FOSSIL [2], and FreeDA [3] are training-free approaches that build prototypes in the visual space according to pre-defined textual categories. CLIP-DINOiser [47] demonstrates that CLIP can be fine-tuned with the supervision of DINO to retain improved localization capabilities. ProxyCLIP [23] proposes a proxy attention module to integrate DINO features with values from the last attention layer of CLIP, thus employing two visual backbones at prediction time. Our method is closely related to this research field since we aim to combine the multimodal understanding capabilities of CLIP with the semantic localization properties of DINOv2. However, in contrast to these methods, we propose to directly map the textual representations from the textual encoder of CLIP to the DINOv2 space, and demonstrate that our approach sets a new state of the art without relying on multiple visual backbones or on external sources of knowledge.
Figure 2: Overview of the training methodology of Talk2DINO. We learn a projection $\psi(\cdot)$ that maps the CLIP textual embeddings to the visual embedding space of DINOv2. Given the dense feature map and the attention maps extracted from DINOv2, we generate $N$ visual embeddings by computing a weighted average of the feature map with each attention map. We then compute the similarity between each visual embedding and the projected text embedding, and apply a selection head function to produce a global similarity score. We use the maximum similarity score to align textual and visual spaces.
3 Proposed Method
3.1 Preliminaries
Task Definition. Open-vocabulary segmentation aims to segment objects of interest defined through natural language at inference time. Let $I\in\mathbb{R}^{H\times W\times 3}$ be an image and $v(I)\in\mathbb{R}^{\frac{H}{P}\times\frac{W}{P}\times D_{v}}$ its dense feature map extracted by a Transformer-based visual backbone with input patch size $P$ and dimensionality of embedding space $D_{v}$. Let $\{T_{j}\}_{j=1,...,M}$ be a set of arbitrary textual categories and $t(T_{j})\in\mathbb{R}^{D_{t}}$ their embeddings extracted by a pre-trained textual backbone. To simplify the notation, in the following, we will refer to $v(I)$ as $v$ and to $t(T_{j})$ as $t_{j}$. Assuming a multimodal setting in which $D_{t}=D_{v}$, we could define the similarity map $\mathcal{S}(I,T_{j})\in\mathbb{R}^{\frac{H}{P}\times\frac{W}{P}}$ for the image $I$ and category $T_{j}$ as the cosine similarity between $t_{j}$ and each spatial entry of $v$. Formally, the similarity map is defined as
$\mathcal{S}(I,T_{j})_{[h,w]}=\frac{v_{[h,w]}\cdot t_{j}^{\intercal}}{||v_{[h,w]}||\,||t_{j}||},$ (1)
where $\cdot_{[h,w]}$ represents indexing over spatial axes. The full resolution similarity map $\hat{\mathcal{S}}(I,T)\in\mathbb{R}^{H\times W}$ is recovered by upsampling $\mathcal{S}(I,T)$ (e.g., via bilinear interpolation). Segmentation masks $\mathcal{M}(I,T_{1},...,T_{M})$ are then derived by assigning pixels to the category with the highest similarity score, i.e.,
$\mathcal{M}(I,T_{1},...,T_{M})_{[h,w]}=\operatorname*{argmax}_{j=1,...,M}{\hat{\mathcal{S}}(I,T_{j})_{[h,w]}}.$ (2)
In order for Eq. 1, and therefore the segmentation from Eq. 2, to work correctly, not only the two $v$ and $t$ spaces should share the same dimensionality, but they should also be constructed so that they also share the same semantics.
CLIP and DINO Duality. Existing vision-language models trained on image-text pairs, such as CLIP [34], can naturally fit the formulation mentioned above, as they provide dense visual and textual embeddings in the same space. However, while CLIP can correctly align global features coming from texts and images (i.e., through the similarities corresponding to CLS tokens), it lacks a precise alignment between the textual feature $t$ and spatial patches $v$.
Conversely, purely visual self-supervised backbones like DINOv2 [33] have shown remarkable semantic and local consistency of spatial embeddings, enabling agnostic image segmentation [39, 38, 45]. These abilities occur naturally in the last attention layer of DINOv2, where the attention maps computed between the CLS token $v^{*}\in\mathbb{R}^{D_{v}}$ and the spatial tokens $v$ align with relevant objects within the image (see Fig. 2). Despite the remarkable results observed on image-only tasks, DINOv2 lacks a solid bridge with natural language, making it impossible to directly replace $v^{*}$ with a text feature $t$ to compute the similarities expressed in Eq. 1.
While DINOv2 and CLIP embedding spaces are traditionally thought as being uncorrelated spaces, we show that the CLIP textual embedding space can be projected into the DINOv2 space through a learnable nonlinear warping followed by an affine transformation.
3.2 Augmenting DINO with Semantics
Warping CLIP Embedding Space. We learn a projection $\psi:\mathbb{R}^{D_{t}}\rightarrow\mathbb{R}^{D_{v}}$ to map textual embeddings $t$ into the space of the visual patch embeddings $v$ of DINOv2, leveraging weak supervision from image-text pairs. We build the projection $\psi$ applied to textual features by composing two affine transformations with a hyperbolic tangent activation, which provides nonlinear warping. Formally,
$\psi(t)=\bm{W}^{\intercal}_{b}\left(\text{tanh}(\bm{W}^{\intercal}_{a}t+b_{a}\right))+b_{b},$ (3)
where $\bm{W}_{a}\in\mathbb{R}^{D_{t}\times D_{v}}$ and $\bm{W}_{b}\in\mathbb{R}^{D_{v}\times D_{v}}$ are learnable projection matrices and $b_{*}$ are learnable bias vectors.
Mapping DINO to the Warped CLIP Space. To learn the nonlinear projection $\psi$, we exploit the intrinsic segmentation capability of DINOv2 to identify the precise spatial subsets of $v$ to which $\psi(t)$ should be aligned with.
Specifically, we first extract the $N$ attention maps $A_{i}\in\mathbb{R}^{\frac{H}{P}\times\frac{W}{P}}$ (one for each of the $i=1,...,N$ heads) which DINOv2 computes between the $v^{*}$ CLS vector and its patch features $v$ from the last layer. One of the key features of DINOv2 is that each $A_{i}$ highlights different semantic regions within the image. For each attention map $A_{i}$, we compute a visual embedding $v^{A_{i}}\in\mathbb{R}^{D_{v}}$ as a weighted average of the dense feature map $v$, emphasizing the spatial areas that $A_{i}$ highlights. We then calculate the cosine similarity between each $v^{A_{i}}$ and the projected text embedding $\psi(t)$, resulting in $N$ similarity scores. Formally, the cosine similarity score between a head and the text embedding is defined as
$\text{sim}(v^{A_{i}},t)$ $=\frac{v^{A_{i}}\cdot\psi(t)^{\intercal}}{||v^{A_{i}}||\,||\psi(t)||},$ (4)
with $\text{with }v^{A_{i}}$ $=\sum_{h,w}v_{[h,w]}\text{softmax}(A_{i})_{[h,w]}.$ (5)
To obtain the most relevant score for alignment, we apply a selection function over the similarity scores obtained for different heads. In particular, we choose the maximum similarity $\max_{i=1,...,N}\text{sim}(A_{i},v,t)$ score across all heads, therefore promoting a robust alignment between textual and visual representations that adapts to the most salient visual features corresponding to the text query.
Training Procedure. To optimize the alignment between text and visual embeddings, we employ the InfoNCE loss, which leverages similarity scores across a batch of image-text pairs. For each pair $(I_{i},T_{i})$, we compute similarity scores between the projected text embedding $\psi(t_{i})$ and the maximally-activated visual embedding $\tilde{v_{i}}$, where $\tilde{v_{i}}$ is the visual embedding derived from the most relevant attention head for the corresponding text $t_{i}$, i.e.,
$\tilde{v}_{i}=v^{A_{j}}_{i}\;|\;j=\operatorname*{argmax}_{k=1,\dots,N}\text{sim}(v_{i}^{A_{k}},t_{i}).$ (6)
Treating the true image-text pair as the positive instance and the remaining pairs within the batch as negatives, this contrastive approach drives the model to increase similarity for matching pairs and decrease it for non-matching pairs. Formally, the InfoNCE loss $\mathcal{L}_{\text{InfoNCE}}$ for a batch of $B$ image-text pairs is defined as
$\mathcal{L}_{\text{InfoNCE}}=-\frac{1}{B}\sum_{i=1}^{B}\log\frac{\exp(\text{sim}(\psi(t_{i}),\tilde{v}_{i}))}{\sum_{j=1}^{B}\exp(\text{sim}(\psi(t_{i}),\tilde{v}_{j}))}.$ (7)
This formulation effectively strengthens alignment by maximizing the similarity for true pairs and minimizing it for mismatched pairs across the batch.
Figure 3: Inference procedure. At the top, we compute the similarity between mapped text embeddings and the DINOv2 patches to produce the initial similarity maps. In the bottom part, we produce a background cleaning map for each class derived from the different DINOv2 attention heads. We obtain the final enhanced similarity map of each category through a convex combination of the similarity and background cleaning maps. The output segmentation then results from the final refinement and thresholding steps.
Inference. The projection learned during the training procedure that warps the CLIP embedding space into the DINOv2 space enables the textual embeddings to be directly comparable with the dense feature embeddings from DINOv2. Hence, at inference time, given an image $I\in\mathbb{R}^{H\times W\times 3}$ and a set of textual arbitrary categories $\{T_{j}\}_{j=1,...,M}$, we can obtain the segmentation masks as defined in Eq. 2 by considering the projected text embeddings $\psi(t_{j})$ in the similarity map computation from Eq. 1.
3.3 Identifying Background Regions
An additional challenge that OVS approaches need to face, especially when tasked with benchmarks like Pascal VOC [15] and COCO Objects [7], is that of identifying “background” regions, i.e. regions that do not belong to the set of categories considered in the benchmark. The standard approach consists in applying a threshold on the similarity or probability score to identify where the model is not certain about the predicted category and classify these locations as background. However, previous works [22, 47, 46] have introduced custom approaches to improve the capabilities of the model in recognizing the background.
Following this line, we propose a background cleaning procedure, depicted in Fig. 3, that is based on the capabilities of the DINOv2 backbone in focusing on coherent areas and highlighting the foreground through the self-attention heads. Specifically, given $N$ attention maps $A_{i}\in\mathbb{R}^{\frac{H}{P}\times\frac{W}{P}}$ and $M$ projected textual embeddings of classes $\psi(t_{j})$, we first compute the average visual embeddings $v^{A_{i}}$ as in Eq. 5. Similarly to the training procedure, we then compute the similarity between each $v^{A_{i}}$ and $\psi(t_{j})$, resulting in a matrix of similarity scores $\mathcal{R}\in\mathbb{R}^{M\times N}$, which is additionally normalized row-wise through a softmax operation. These scores represent how much each self-attention head is related to each textual category. Formally, $\mathcal{R}$ is defined as
$\mathcal{R}$ $=[\mathcal{R}_{1},...,\mathcal{R}_{j},...,\mathcal{R}_{M}]^{\intercal},\text{ with }$ (8)
$\mathcal{R}_{j}$ $=\text{softmax}\left(\text{sim}(v^{A_{1}},\psi(t_{j})),...,\text{sim}(v^{A_{N}},\psi(t_{j}))\right).$
Then, for each category $T_{j}$ we compute its average attention map $\mathcal{F}_{j}\in\mathbb{R}^{\frac{H}{P}\times\frac{W}{P}}$ as
$\mathcal{F}_{j}(A_{i},\mathcal{R}_{ij})=\sum_{i=1}^{N}\mathcal{R}_{ij}A_{i},$ (9)
and normalize $\mathcal{F}$ by applying a softmax normalization over both spatial axes, and linearly re-projecting its values in the range $[\min_{j,h,w}\mathcal{S}(I,T_{j})_{[h,w]},\max_{j,h,w}\mathcal{S}(I,T_{j})_{[h,w]}]$, where $\mathcal{S}(\cdot)$ is the similarity map defined in Eq. 1. We exploit the resulting normalized average attention per category to shape the similarity map by activating the foreground region and deactivating the background. The resulting shaped similarity map $\overline{\mathcal{S}}\in\mathbb{R}^{H\times W}$ is defined as
$\overline{\mathcal{S}}(I,T_{j})_{[h,w]}=\lambda\mathcal{S}(I,T_{j})_{[h,w]}+(1-\lambda)\mathcal{F}_{j,[h,w]},$ (10)
where $\lambda$ is a hyperparameter representing the relevance of the background shaping in computing the segmentation masks. The segmentation mask for the background region is then simply identified as the collection of pixels for which the shaped similarity map is lower than a threshold across all semantic categories.
ViT-Base (mIoU) ViT-Large (mIoU)
Model Visual Backbone Frozen VOC Context Stuff Cityscapes ADE Avg VOC Context Stuff Cityscapes ADE Avg
without Mask Refinement
GroupViT [49] Custom ViT 79.7 23.4 15.3 11.1 9.2 27.7 - - - - - -
ReCo [37] CLIP 57.7 22.3 14.8 21.1 11.2 25.4 - - - - - -
TCL [9] CLIP 77.5 30.3 19.6 23.1 14.9 33.1 - - - - - -
SILC [32] Custom ViT 77.5 31.6 20.8 26.9 19.3 35.2 - - - - - -
MaskCLIP [58] CLIP 72.1 25.3 15.1 11.2 9.0 26.5 29.4 12.4 8.8 11.5 7.2 13.9
CLIP-DIY [46] CLIP+DINO 79.7 19.8 13.3 11.6 9.9 26.9 - - - - - -
OVDiff [22] CLIP+DINO+SD 80.9 32.9 20.3 23.4 14.9 34.5 - - - - - -
SCLIP [42] CLIP 80.4 34.2 22.4 32.2 16.1 37.1 70.6 25.2 17.6 21.3 10.9 29.1
CLIP-DINOiser [47] CLIP 80.9 31.7 24.6 35.9 20.0 38.6 - - - - - -
NACLIP [19] CLIP 79.7 35.2 23.3 35.5 17.4 38.2 78.7 32.1 21.4 31.4 17.3 36.2
ProxyCLIP [23] CLIP+DINOv2 83.0 37.2 25.4 33.9 19.7 39.8 85.2 36.2 24.6 35.2 21.6 40.6
ProxyCLIP [23] CLIP+DINO 80.3 39.1 26.5 38.1 20.2 40.8 83.2 37.7 25.6 40.1 22.6 41.8
Talk2DINO (Ours) DINOv2 85.3 40.5 27.9 38.2 21.8 42.7 86.8 39.3 27.1 37.0 21.4 42.3
with Mask Refinement
GroupViT [49] Custom ViT 81.5 23.8 15.4 11.6 9.4 28.3 - - - - - -
ReCo [37] CLIP 62.4 24.7 16.3 22.8 12.4 27.7 - - - - - -
TCL [9] CLIP 83.2 33.9 22.4 24.0 17.1 36.1 - - - - - -
MaskCLIP [58] CLIP 74.9 26.4 16.4 12.6 9.8 28.0 32.7 14.6 10.5 13.2 8.5 15.9
FOSSIL [2] DINOv2 - - - - - - - 35.8 24.8 23.2 18.8 -
SCLIP [42] CLIP 83.5 36.1 23.9 34.1 17.8 39.1 76.3 27.4 18.7 23.9 11.8 31.6
NACLIP [19] CLIP 83.0 38.4 25.7 38.3 19.1 40.9 84.5 36.4 24.6 37.1 19.6 40.4
FreeDA [3] DINOv2 75.6 37.6 25.2 34.4 20.7 38.7 70.2 39.0 26.9 33.2 19.5 37.8
FreeDA [3] CLIP+DINOv2 85.6 43.1 27.8 36.7 22.4 43.1 87.9 43.5 28.8 36.7 23.2 44.0
Talk2DINO (Ours) DINOv2 87.0 43.5 30.3 40.8 23.5 45.0 88.7 43.2 30.0 39.3 23.4 44.9
Table 1: Comparison with unsupervised OVS models on Pascal VOC [15], Pascal Context [30], COCO Stuff [7], Cityscapes [11], and ADE20K [56, 57]. For each method, we specify the visual backbone used, along with whether it is frozen or fine-tuned.
4 Experiments
4.1 Experimental Setup
Datasets. We evaluate our approach on eight widely-used semantic segmentation benchmarks, which we categorize based on the inclusion of a background class. Specifically, we conduct experiments on the validation sets of Pascal VOC 2012 [15], Pascal Context [30], COCO Stuff [7], Cityscapes [11], and ADE20K [56, 57], that contain 20, 59, 171, 150, and 19 semantic categories respectively and do not include the “background” class. We report additional experiments on the COCO Objects dataset [7], which consists of 80 different foreground object classes, and on modified versions of Pascal VOC 2012 and Pascal Context in which the “background” category is, instead, included.
Implementation Details. For the main experiments, we employ DINOv2 ViT-B/14 as the base model and DINOv2 ViT-L/14 as the large model, both with the CLIP ViT-B/16 text encoder. We use the DINOv2 variant with registers [12] since our method benefits from the removal of artifacts in self-attention maps [^1] A detailed analysis on the impact of using DINOv2 with registers is reported in the supplementary material.. We resize the input image to $518\times 518$ to match the original training resolution of DINOv2, obtaining $37\times 37$ patches. We train the model with the Adam optimizer, a batch size of $128$, and a learning rate of $1\times 10^{-4}$ for a total of 100 epochs on the COCO Captions 2014 training split [27], composed of around 80k images.
Following previous works [9, 22, 19], we optionally employ a mask refinement stage to counteract any inaccuracies in the final masks. In particular, we adopt Pixel-Adaptive Mask Refinement (PAMR) [1], an iterative post-refinement method aimed at enhancing the fidelity of the similarities to the visual characteristics of the image. In our experiments, we use $\lambda$ equal to $\nicefrac{{5}}{{6}}$ for background cleaning, a threshold of 0.55 on the similarity score to determine which pixels belong to the “background” category, and, when using mask refinement, employ PAMR with 10 iterations.
Evaluation Protocol. We follow the standard evaluation protocol for unsupervised OVS [9], where prior access to the target data before evaluation is not allowed, and use the default class names provided by the MMSegmentation toolbox. All models are evaluated using mean Intersection-over-Union (mIoU) on all the classes of each dataset. The images are resized to have a shorter side of 448, using a sliding window approach with a stride of 224 pixels.
4.2 Comparison with the State of the Art
We compare Talk2DINO with previous state-of-the-art approaches for unsupervised OVS on the five benchmarks that do not include the “background” category and the three benchmarks with the “background” category. We consider as competitors: (i) prototype-based approaches, such as ReCo [37], OVDiff [22], FOSSIL [2], and FreeDA [3], which aim to create visual prototypes associated with the textual categories, (ii) CLIP adaptations, as MaskCLIP [58], CLIP-DIY [46], SCLIP [42], and NACLIP [19], which propose architectural modifications to enhance the localization properties of CLIP, (iii) methods trained on large-scale sets of image-captions pairs with loss functions designed to force the segmentation capabilities to emerge, like GroupViT [49], ViewCo [36], TCL [9], OVSegmentor [50], SegCLIP [28], and SILC [32], and (iv) methods that aim to combine the complementary properties of CLIP and DINO, like CLIP-DINOiser [47] and ProxyCLIP [23].
Table 1 reports the results on the five benchmarks without background (i.e., Pascal VOC, Pascal Context, COCOStuff, Cityscapes, and ADE). Specifically, we report the performance of both the base and large configurations of both Talk2DINO and the competitors, according to their definitions in the original papers. Moreover, we divide the table into two sections depending on whether a mask refinement technique is employed. Specifically, FreeDA exploits Felzenszwalb’s superpixel algorithm [17], FOSSIL introduces a custom region proposal algorithm named OpenCut, while all the other methods refine their masks with PAMR. As it can be observed, our approach achieves the best average mIoU on all the configurations and presents a consistent improvement compared to the considered competitors, with and without the mask refinement, across all datasets.
mIoU
Model Visual Backbone Frozen VOC Context Object
ViT-Base
GroupViT [49] Custom ViT 51.1 19.0 27.9
ViewCo [36] Custom ViT 52.4 23.0 23.5
OVSegmentor [50] DINO 53.8 20.4 25.1
SegCLIP [28] CLIP 52.6 24.7 26.5
TCL [9] CLIP 55.0 30.4 31.6
MaskCLIP [58] CLIP 37.2 22.6 18.9
CLIP-DIY [46] CLIP 59.9 19.7 31.0
OVDiff [22] CLIP+DINO+SD 68.4 31.2 36.2
SCLIP [42] CLIP 61.7 31.5 32.1
CLIP-DINOiser [47] CLIP 62.1 32.4 34.8
NACLIP [19] CLIP 64.1 35.0 36.2
FreeDA [3] DINOv2 52.9 35.5 25.9
FreeDA [3] CLIP+DINOv2 55.4 37.9 36.2
ProxyCLIP [23] CLIP+DINOv2 58.6 33.8 37.4
ProxyCLIP [23] CLIP+DINO 61.3 35.3 37.5
Talk2DINO (Ours) DINOv2 65.9 38.4 44.6
ViT-Large
NACLIP [19] CLIP 57.9 36.4 34.6
FreeDA [3] DINOv2 50.3 34.3 26.7
FreeDA [3] CLIP+DINOv2 55.4 38.3 37.4
ProxyCLIP [23] CLIP+DINOv2 56.6 33.0 36.7
ProxyCLIP [23] CLIP+DINO 60.6 34.5 39.2
Talk2DINO (Ours) DINOv2 66.6 37.7 43.0
Table 2: Comparison with unsupervised OVS models on Pascal VOC [15], Pascal Context [30], and COCO Object [7], considering the background category. For each method, the visual backbone used is specified, along with whether it is frozen or fine-tuned. We report all methods in their best configuration.
Table 2, instead, shows the results on the three benchmarks that require assigning the “background” category to pixels that do not belong to the set of provided categories. Pascal VOC and COCO Objects present only foreground categories, also referred to as “things” in the literature, while Pascal Context presents both categories from foreground and background, also mentioned as “stuff”. Hence, following CLIP-DINOiser [47], we report the performance with the background cleaning procedure only on Pascal VOC and COCO Objects. From the results, it can be seen that Talk2DINO achieves the best or second-best result on all considered configurations. We also notice that OVDiff builds specific negative prototypes for each category to enhance its representation of the background.
The most straightforward comparison is the one with FreeDA without global similarity (i.e., with DINOv2 only as visual backbone). Indeed, this approach builds a bridge between DINOv2 and the CLIP text encoder by retrieving from a collection of visual and textual embedding pairs and by building visual prototypes for each textual category. The significant improvement achieved by Talk2DINO demonstrates that training a direct projection from the CLIP text encoder to DINOv2 leads to a more accurate bridge between the two embedding spaces without the overhead in computation and memory provided by the retrieval procedure.
4.3 Ablation Studies and Analyses
Choosing Different Visual Backbones. In Table 3 we show the performance of our approach when varying the visual backbone and the size of the employed ViT architecture. In these experiments, we keep the original resolution of the models during the training procedure and follow the same downsampling protocol employed for Talk2DINO at test time. We observe that backbones that differ from DINOv2 present unsatisfactory results and can not be aligned to the CLIP textual encoder with a learnable mapping. In particular, while DINO achieves the second best performance on average, Talk2DINO heavily benefits from the strong semantic representation of the dense features of DINOv2 and on the capabilities of its self-attention heads in highlighting coherent regions of the image – properties which are not reflected in other visual backbones (see supplementary materials for detailed analysis). Additionally, we notice that our approach demonstrates consistent and high performance across different ViT sizes, with strong results even when using the ViT-Small backbone.
mIoU
Visual Backbone VOC Context Stuff Cityscapes ADE
DINO ViT-S 25.8 12.6 7.0 15.2 4.9
DINOv2 ViT-S 86.8 35.3 24.6 28.3 17.0
MAE ViT-B 5.3 2.7 1.4 4.7 0.9
CLIP ViT-B 47.3 9.7 9.8 3.3 2.4
DINO ViT-B 22.0 12.3 6.9 16.4 4.7
DINOv2 ViT-B 85.3 40.5 27.9 38.2 21.8
MAE ViT-L 6.7 2.7 1.4 4.6 0.9
CLIP ViT-L 16.3 4.9 7.4 1.0 2.0
DINOv2 ViT-L 84.7 39.4 27.2 36.3 21.2
Table 3: Ablation study results using different visual backbones and with different sizes of the ViT architecture.
mIoU
VOC Context Stuff Cityscapes ADE
Effect of Projection
Single FC (text only) 84.5 38.5 26.9 35.3 20.1
Our mapping (both vision and text) 56.8 23.9 16.8 21.9 12.6
Our mapping (vision only) 77.1 30.6 20.9 22.5 15.4
Our mapping (text only) 85.3 40.5 27.9 38.2 21.8
Effect of Self-Attention Selection and Aggregation
CLS only (without self-attention) 85.4 30.4 22.5 22.9 16.4
Standard average 89.2 38.6 26.0 34.8 20.4
CLS-weighted average 87.4 34.9 22.8 29.5 16.9
CLS similarity-weighted sampling 88.2 32.4 22.7 29.9 18.8
Max CLS similarity 85.3 40.5 27.9 38.2 21.8
Table 4: Ablation study evaluating the impact of the core components of the proposed architecture on the final performance. We report the results using the base model of DINOv2.
Image Ground-truth FreeDA [3] ProxyCLIP [23] CLIP-DINOiser [47] Talk2DINO (Ours)
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 4: Qualitative results of Talk2DINO in comparison with FreeDA [3], ProxyCLIP [23], and CLIP-DINOiser [47].
Impact of the Proposed Components. Table 4 reports the results of Talk2DINO evaluating the impact of its core components on the overall performance. Specifically, in the first section of the table, we analyze the effect of the adopted projection $\psi$. Noticeably, replacing it with a linear projection leads to a slight drop in performance. The good performance obtained by a linear transform evidences how the DINOv2 and CLIP spaces are intrinsically compatible, as the former can be obtained through an affine transform of the latter without losing too much information. Interestingly, applying the proposed projection on top of DINOv2 or applying two projections on both spaces leads to significantly lower performance, confirming the appropriateness of the proposed approach.
In the second section of the table, we instead study the effect of the selection and aggregation strategy of self-attention heads $A_{i},i=1,...,N$ during training. In particular, we test aligning (i) directly the visual CLS token to the textual CLS token, (ii) the visual embedding from the standard average self-attention, (iii) the weighted mean of the head embeddings $v_{A_{i}}$, where the weights are given by their softmaxed similarity with the textual CLS token, (iv) a strategy in which we sample a single head embedding, where the sampling probability is given by the softmaxed similarity with the CLS token, and (v) our adopted solution in which we select the head embedding which is the most similar to the textual CLS token. The results show that only on the Pascal VOC dataset the embedding from the standard average self-attention presents improved performance. On all other benchmarks, our approach proves most effective, further validating the robustness of our selection method.
mIoU
Mask Refinement VOC Context Object
without background cleaning 57.3 35.7 36.5
with background cleaning 61.1 35.3 40.4
without background cleaning 61.2 38.4 39.7
with background cleaning 65.9 37.6 44.6
Table 5: Ablation study comparing results with and without the background cleaning procedure. We report the results using the base model of DINOv2.
Effect of Background Cleaning. Table 5 shows how the performance is affected by the background cleaning mechanism and by the usage of PAMR for mask refinement. It can be observed that, according to Sec. 4.2, background cleaning has a positive impact on Pascal VOC and COCO Object, which contain only foreground categories, while slightly degrades the performance on Pascal Context, which however contains categories that are usually part of the background of an image (e.g., sky, road, and wall). Further, it can be noticed that the effectiveness of the proposed background cleaning procedure is confirmed both with and without mask refinement.
Qualitative Results. Fig. 4 depicts qualitative segmentation results, in which we highlight the segmentation capabilities of Talk2DINO along with other state-of-the-art models [3, 23, 47]. We show two images from Pascal VOC and COCO Object, in which it can also be appreciated how the background cleaning procedure leads to high-quality masks and localization. Moreover, we show two images from Cityscapes and Pascal Context where Talk2DINO effectively segments small objects in the scene, such as the traffic signs, and “stuff” categories, such as sky and road.
5 Conclusion
In this paper we introduced Talk2DINO, a novel approach for open-vocabulary segmentation that bridges the spatially detailed embeddings of the DINOv2 self-supervised vision backbone with the highly semantic CLIP’s text embeddings. Our method achieves fine-grained alignment between textual concepts and visual patches without the need for extensive fine-tuning of the backbone networks, only leveraging DINOv2’s self-attention maps and a lightweight language-to-vision mapping layer. Talk2DINO achieves state-of-the-art results in standard OVS benchmarks, outperforming previous prototype-based, CLIP adaptation, and contrastive learning approaches. Our study highlights the potential of integrating vision-only and multimodal models, suggesting broader applications in fine-grained and cross-modal tasks.
Acknowledgments
We acknowledge the CINECA award under the ISCRA initiative, for the availability of high-performance computing resources and support. This work has been conducted under a research grant co-funded by Leonardo S.p.A., and supported by the PNRRM4C2 project “FAIR - Future Artificial Intelligence Research”, by the PRIN 2022-PNRR project “MUCES” (CUP E53D23016290001 and B53D23026090001), and by the PNRR project “ITSERR - Italian Strengthening of Esfri RI Resilience” (CUP B53C22001770006), all funded by the European Union - NextGenerationEU, and by the SUN XR project funded by the Horizon Europe Research & Innovation Programme (GA n. 101092612).
Supplementary Material
Figure 5: Performance vs. Parameter Count. The y-axis denotes the obtained mIoU averaged over all the five benchmarks reported in the main paper. Dashed lines connect methods tested with multiple backbone sizes; these are labeled S/B/L for Small/Base/Large ViT models. Talk2DINO offers the best trade-off.
Appendix A Additional Experiments and Analyses
Analysis of Model Parameters. Fig. 5 reports a comparison of the relationship between the performance, in terms of average mIoU, and the number of parameters of the models. As it can be observed, Talk2DINO presents a lower number of parameters than the recent competitors FreeDA [3] and ProxyCLIP [23], along with an improved average mIoU. Models with a comparable number of parameters, such as TCL [9], GroupViT [49], and MaskCLIP [58], exhibit a lower performance compared to Talk2DINO. Finally, it shall be noted that models such as FreeDA, ReCo [37], and OVDiff [22] require maintaining external sources of knowledge, which increases memory consumption.
Role of DINO Registers. The main configuration of Talk2DINO, with both the base and large sizes, leverages the variant of DINOv2 with registers. Registers are a recently proposed mechanism (see Darcet et al. [12]) to mitigate the presence of artifacts in the feature maps of ViT-based backbones [14]. Artifacts are tokens that exhibit a significantly higher norm with respect to the other tokens and retain less information about their original position in the image. This phenomenon has been especially observed after long training and is more pronounced in backbones with larger sizes. In Fig. 8 we depict, on the first row, the average self-attentions between the CLS and the other tokens for the ViT-S, ViT-B, and ViT-L architectures with and without registers, while in the following rows, we show the various self-attention heads for each backbone. It can be observed that in the ViT-S the artifacts are not present, and the average self-attention between the model with and without the registers is nearly identical. Instead, the ViT-B exhibits artifacts in the top left corner, resulting in an average self-attention that is especially focused on that portion and less active on the people. This side effect is even more noticeable with the ViT-L, for which the artifact is the only visible token in the average self-attention.
In Table 6, we compare the performance achieved when employing the DINOv2 visual backbone with and without registers. The results show a downgrade in performance without the registers that is directly related to the presence of the artifacts in the self-attentions, according to the observations from Fig. 8. Indeed, the largest difference in performance is measured in the ViT-L architecture, while in the ViT-S case, the backbone without registers performs better on four benchmarks out of five. The phenomenon of the artifacts represents a threat to our approach because it limits the mechanism of selection of the self-attention head most similar to the caption, forcing the alignment between the unwanted artifact tokens and the caption.
Effect of Training CLIP Last Layer. Table 7 reports a comparison between Talk2DINO when training only the $\psi(t)$ projection as proposed in the main paper and when instead unfreezing the last layer of CLIP [34]. Despite this experiment exhibiting a small performance gap between the two configurations, unfreezing the last layer of CLIP, interestingly, leads to worse results. This outcome highlights that the textual representations provided by CLIP, which have been pre-trained to match their visual counterpart, if trained inside a different pipeline, can be harmed and can lose part of their capabilities in multimodal understanding.
Choosing Different Visual Backbones. In Table 3 of the main paper, we reported the performance of our approach applied to different visual backbones. The results demonstrated that our training pipeline is especially suitable for DINOv2, while it leads to unsatisfactory performance on DINO, MAE, and CLIP. In this paragraph, we further explore, from a qualitative perspective, the limitations of these probed visual backbones. Fig. 9 showcases the average self-attentions between the CLS and the other tokens on the first row, breaking down the contributions from the various self-attention heads in the successive rows. We can notice that the self-attention heads of CLIP present a noise that, similar to what we observed for DINOv2 without registers, represents a limitation for our proposed training procedure. On the contrary, the self-attention heads of DINO and MAE are clean and highlight homogeneous regions of the image. In these cases, the performance gap with DINOv2 can be related to the insufficient semantic richness of the dense features of these backbones. These insights are also confirmed by the ablation study reported in FreeDA [3], in which a similar performance gap has been observed when using the CLIP and DINO backbones instead of DINOv2.
mIoU
Visual Backbone VOC Context Stuff Cityscapes ADE
DINOv2 (without registers) ViT-S 83.7 38.3 25.7 32.9 19.8
DINOv2 (with registers) ViT-S 86.8 35.3 24.6 28.3 17.0
DINOv2 (without registers) ViT-B 68.3 28.5 20.5 20.9 13.6
DINOv2 (with registers) ViT-B 85.3 40.5 27.9 38.2 21.8
DINOv2 (without registers) ViT-L 58.6 21.2 15.8 14.5 8.9
DINOv2 (with registers) ViT-L 84.7 39.4 27.2 36.3 21.2
Table 6: Performance comparison of DINOv2 backbones with and without registers.
Using CLIP Text Tokens. In Table 7, we report the results of utilizing the dense output of the CLIP text encoder instead of its CLS token for alignment. While our primary experiments align the CLS token with the best attention map embedding to target the patches most relevant to the text, we also explore aligning individual text tokens to the best attention map embeddings. This approach is motivated by the hypothesis that each word in the text might correspond to a distinct region in the image. During inference, since we perform the alignment on individual text tokens rather than the CLS token, we average the text tokens to calculate similarity with the visual patches. However, this method yields inferior results compared to using the CLS token. We then refine this approach by aligning only a subset of text tokens selected using nucleus sampling ($\alpha=0.6$) to filter out potentially irrelevant words, such as stop words. Despite this effort, performance does not improve.
These observations suggest that the global objective of CLIP’s training, similar to its effect on visual patch embeddings, may not endow text tokens with strong local properties that accurately reflect the specific word each embedding represents. This limitation likely contributes to the noisiness of such alignments. Additionally, we evaluate the use of the average of CLIP text tokens in both training and inference as an alternative to the CLS token. While this approach slightly improves over aligning individual tokens, it still underperforms compared to the CLS token, indicating that it encapsulates the most useful and less noisy information for alignment with DINOv2 patches.
mIoU
VOC Context Stuff Cityscapes ADE
Effect of Training CLIP Last Layer
Trained 83.0 36.9 25.8 30.3 19.0
Frozen 85.3 40.5 27.9 38.2 21.8
Effect of Text Token Selection
Average text tokens 83.2 37.8 25.5 32.7 19.6
Text token to best self-attn map 84.2 34.8 24.2 32.9 18.7
Text token to best self-attn map (NS) 81.7 33.4 23.8 29.1 18.8
CLS token only 85.3 40.5 27.9 38.2 21.8
Table 7: Ablation study on the effect of training the last layer of CLIP and text token selection strategies.
[…]
R@1 $\uparrow$ R@5 $\uparrow$ R@10 $\uparrow$ Median $\downarrow$ Mean $\downarrow$ R@1 $\uparrow$ R@5 $\uparrow$ R@10 $\uparrow$ Median $\downarrow$ Mean $\downarrow$
ViT-Base
CLIP 41.3 65.8 76.3 2 13.4 22.6 44.1 54.9 8 52.5
Talk2DINO 32.9 59.2 71.1 3 14.6 12.2 34.3 48.7 11 34.2
   + Custom Alignment 39.4 67.7 78.2 2 10.7 26.2 53.0 66.4 5 21.0
ViT-Large
CLIP 45.4 71.1 79.2 2 11.0 26.5 48.7 59.0 6 44.2
Talk2DINO 26.5 53.7 65.6 5 18.8 12.7 33.7 47.8 11 43.1
   + Custom Alignment 37.9 64.7 75.1 3 13.1 24.4 50.1 63.2 5 27.8
Table 8: Retrieval performance on the COCO Captions test set.
Appendix B Image-Text Matching Results
While Talk2DINO is primarily designed for OVS, we also assess its performance on image-text retrieval to evaluate its capabilities in global image understanding. For this task, we adopt the same text encoding approach used in segmentation, projecting the CLIP text embedding. The global image representation is derived by averaging the embeddings computed from each DINOv2 attention map. Specifically, for each attention map $A_{i}$, we calculate a visual embedding $v^{A_{i}}\in\mathbb{R}^{D_{v}}$ as the weighted average of the dense feature map $v$. The final global image representation is then obtained by taking the mean of all $v^{A_{i}}$ embeddings.
In Table 8, we assess the retrieval performance on the COCO Captions test set [27] using both ViT-B and ViT-L configurations. While Talk2DINO generally performs slightly below CLIP across most metrics, it demonstrates a notable advantage in the mean rank for the text-to-image retrieval task. This result underscores the ability of Talk2DINO to better address extreme failures compared to CLIP, indicating improved robustness in handling challenging or outlier queries. In addition to computing text-image similarities using cosine similarity between a global text token and a global image token, we experiment with a similarity function that mirrors the one used during training. Specifically, instead of representing the image with the mean of the $v_{A_{i}}$ embeddings and calculating similarity as the cosine similarity between this representation and the text encoding, we represent the image using all $v_{A_{i}}$ embeddings. We compute the similarity as $\max_{i=1,\dots,N}\text{sim}(A_{i},v,t)$, taking the maximum similarity score across all heads. This alternative similarity function leads to significant performance improvements, allowing Talk2DINO to surpass CLIP on several metrics. This enhancement is likely due to the ability of the model to evaluate captions at a finer granularity. Captions often describe multiple aspects of an image, including both foreground and background elements. By individually examining different regions of the image as detected by distinct attention heads, the model can assign more precise scores, ultimately boosting retrieval accuracy.
Appendix C Activation Map Visualizations
In Fig. 6, we show the distribution of attention heads selected for alignment with the text input during the final epoch of training. The results indicate that certain heads, particularly heads 0, 4 (by far the most frequently selected), 8, and 12, are more often aligned with the text than others. However, aside from these, the remaining heads are relatively evenly distributed. These findings are noteworthy because they suggest that some heads specialize in capturing features that align more closely with the input caption, while all heads contribute meaningfully during training. Notably, no head shows a negligible activation frequency, highlighting the importance of the entire set of attention heads in the alignment process.
Figure 6: Percentage of times each attention head of ViT-B backbone is selected for alignment to textual embeddings on the final epoch of training. The dashed line denotes uniform distribution.
Fig. 7 presents examples from the training set, showcasing images paired with their corresponding captions and the attention maps selected for alignment. Despite describing the same scene, variations in the captions lead the alignment procedure to focus on different regions of the image. For instance, in the first row, a caption mentioning only the food on the table excludes the dog from the focus, while captions that also reference the dog include it in the relevant patches.
Head 3: A cut up sandwich on top of a wooden table.
Head 8: a dog is stepping up onto the table to eat a plate of food
Head 11: A small dog is sitting at the kitchen table in front of a sandwich.
Head 12: A dog in a shirt sits at a table with a plate of food in front of the dog.
Head 15: A Chihuahua sits at a table next to a sandwich.
Head 0 A baseball games is underway with someone on second base as the pitcher throws the ball.
Head 2: A baseball game from a distance during a pitch.
Head 6: A professional baseball game between to MLB teams
Head 11: baseball players on a baseball field during a professional game
Head 12: Baseball players are ready for the next pitch.
Head 2: A white round table filled with some assorted treats.
Head 5: a dessert tray with dougnuts, cupcakes and muffins
Head 6: A white plate topped with chocolates, donuts and other goodies.
Head 9: A number of pastry items on a table
Head 15: Donuts and other goodies on a table
Head 0: A surfer in a bright green suit expertly takes a wave.
Head 2: a surfer is out at sea riding a wave
Head 4: A surfer turns to catch the oncoming wave.
Head 10: A man who is on a surfboard riding a wave.
Head 14: A man surfing in strong waves in the sea
Head 1: A zebra is standing in an enclosure in a zoo.
Head 3: A zebra is in an enclosure by a fence.
Head 7: A zebra standing near another zebra in a enclosure.
Head 9: A zebra standing in the dirt in an enclosure.
Head 10: A picture of zebra in a enclosed area.
Figure 7: Sample images from the training set paired with their corresponding captions and the attention maps selected for alignment during the last epoch of training.
Appendix D Additional Qualitative Results
Effect of Background Cleaning. In Fig. 10 we showcase a set of qualitative results in which we highlight the advantages of using the proposed background cleaning procedure with respect to directly thresholding the similarities with the provided categories to detect the background. In particular, the first two rows show four qualitatives on images from COCO Object and the last two rows from Pascal VOC. These results demonstrate that background cleaning removes the noise in the background from the image and improves the fitting of the masks on the foreground objects. These findings are reflected in the results reported in Table 5 of the main paper.
In-the-Wild Qualitative Examples. Fig. 11 depicts a few examples of “in-the-wild” segmentation, obtained by providing to Talk2DINO some images retrieved from the web and asking it to segment uncommon categories, such as “pikachu”, “millennium falcon”, and “westminster abbey”, and free-form text, like “golden retriever puppy”. On the left, we show three examples in which we task the model with also finding the background, while exploiting the background cleaning procedure, and, on the right, three examples in which the model has to assign a provided category to each pixel. The high quality of the resulting masks demonstrates the efficacy of our approach, even on out-of-domain images with respect to the standard benchmarks. From these examples, we can appreciate the model’s capabilities in combining the knowledge from CLIP with the semantic localization of DINOv2 on unconventional concepts, such as fictional character names and proper nouns of buildings and monuments.
Comparison with State-of-the-Art Methods. Finally, in Fig. 12 we report a set of qualitative results on the five datasets used for the evaluation of the models, in addition to the qualitative depicted in Fig. 4 of the main paper. We compare the segmentation masks of Talk2DINO with the ones of FreeDA [3], ProxyCLIP [23], and CLIP-DINOiser [47], which represent our main competitors. In particular, we report a pair of images from Pascal VOC with background and eight pairs of images from Pascal Context, COCO Stuff, Cityscapes, and ADE20K, without background. As it can be seen, these qualitative results further highlight the impressive segmentation capabilities of Talk2DINO with both background and foreground categories.
Image ViT-S ViT-S Reg. ViT-B ViT-B Reg. ViT-L ViT-L Reg.
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
ViT-S
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
ViT-S Reg.
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
ViT-B
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
ViT-B Reg.
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
ViT-L
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
ViT-L Reg.
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 8: Comparison of DINOv2 with and without registers across different visual backbones (ViT-S, ViT-B, and ViT-L). The results highlight how the ViT-B and ViT-L backbones without registers exhibit artifacts that introduce noise during the alignment process.
Image DINOv2 DINO MAE CLIP
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
DINOv2
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
DINO
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
MAE
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
CLIP
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption 2 Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 9: Self-attention activations of different visual backbones (i.e., DINOv2, DINO, MAE, CLIP).
with without with without
Image Ground-truth Bkg. Cleaning Bkg. Cleaning Image Ground-truth Bkg. Cleaning Bkg. Cleaning
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 10: Qualitative results obtained with and without the proposed background cleaning strategy, on COCO Object and Pascal VOC.
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
moai dinosaur man westminster abbey golden retriever puppy grass leaves pikachu traffic sign forest street millennium falcon planet space
Figure 11: ”In-the-wild” segmentation results obtained by prompting Talk2DINO with uncommon textual categories on images retrieved from the web.
Image Ground-truth FreeDA [3] ProxyCLIP [23] CLIP-DINOiser [47] Talk2DINO (Ours)
Pascal VOC Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Pascal Context Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
COCO Stuff Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Cityscapes Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
ADE20K Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 12: Additional qualitative results of Talk2DINO in comparison with FreeDA [3], ProxyCLIP [23], and CLIP-DINOiser [47].